A gas vowel from an explosion under water osscilates with a period T and depends on pressure and density of water and energy . derive the formula.

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Defining firstly the parameters below:
P = weight of water 
d = thickness of water 
E = vitality of water 
Day and age is specifically propor tional to the weight 
Tα P - (1.) 
As we likewise realize that 
T straightforwardly corresponding to the thickness of water 
T α d - (2.) 
T is straightforwardly corresponding to vitality of water 
Tα E - (3.) 
Join each of the three 
T α PdE 
T = k PdE 
K = dimensionless steady 
T = k PadbEc 
Measurement of T = [T] 
Measurement of P =[ML-1T-2] 
Measurement of d = [ML-3T0] 
Measurement of E = [ML2T-2] 
[T] = [ML-3T0]b[ ML-1T-2]a [ML2T-2]c 
At that point 
a + b+ c= 0 
- 3b-a+2c = 0 
- 2a-2c = 1 
a = - 5/6 
b =1/2 
c = 1/3 
So the equation is 
T = k P-5/6d1/2 E1/3 
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)